# CogAD: Cognitive-Hierarchy Guided End-to-End Autonomous Driving

### v2

@@ 1 Introduction
Figure 1: Human drivers perceive holistically (c) before focusing on critical objects (d), and plan hierarchically from intent to trajectory (b).
@@
In the perception pattern, we argue that the perception of human drivers is hierarchical. Cognitive psychology supports the pre-attentive two-stage theory of human visual perception [27], which posits that individuals first perform global scanning of the scene to identify contextual or spatial cues, followed by selective attention allocation to critical local objects for detailed analysis. This mechanism aligns with neurobiological evidence [13] demonstrating that early visual areas prioritize scene-wide features, while higher-order regions refine task-specific processing through context-dependent modulation. In terms of driving environmental perception, human drivers typically initiate their environmental perception by establishing holistic scene comprehension of the traffic scenario, encompassing all contextual elements as illustrated in Figure 1 (c), followed by selective attention prioritization towards critical dynamic objects such as proximal vehicles, pedestrians, and lane markings, as illustrated in Figure 1 (d).
In the planning pattern, we argue that the planning of human drivers is hierarchical. The BDI (Belief-Desire-Intention) cognitive model [2, 41] decomposes human action planning into hierarchically organized processes, distinguishing between high-level intentional states (e.g., goal setting) and low-level behavioral executions (e.g., motor sequence implementation). In terms of driving planning, human drivers typically follow a hierarchical planning process: intentional planning at the higher level (e.g., lane change) followed by trajectory planning at the lower level (e.g., generating a specific trajectory). For instance, drivers first establish a global intent, which then constrains the generation of local trajectories, as illustrated in Figure 1 (b).
From the perspective of hierarchical cognitive psychology, recent methods like [55, 24, 19] fail to incorporate hierarchical perception, while others like [47, 23, 33] lack hierarchical planning, as detailed in Section 2. These analyses highlight a critical gap in aligning E2E-AD models with human drivers’ cognitive principles. To bridge this gap, we propose CogAD, an end-to-end autonomous driving model that emulates human hierarchical cognition. For hierarchical perception, we design a sequential interaction mechanism where the ego-vehicle first processes global BEV features to capture environmental context, then focuses on critical instance-level elements. For hierarchical planning, CogAD first plans high-level driving intents, then generates corresponding low-level trajectories, mirroring human multi-level planning.
@@ 2 Related Work
Significant progress has been made in end-to-end autonomous driving (E2E-AD) approaches, pioneered by prior works such as [1, 10, 38]. We analyze the E2E-AD approaches from hierarchical perception and hierarchical planning perspectives.
Hierarchical Perception.
@@ 3 Method
Overview.
As depicted in Figure 2, the proposed CogAD implements a unified multi-task architecture that integrates perception, prediction, and planning in a coordinated manner. To comprehensively capture environmental information, we design a hierarchical feature interaction framework. First, we establish interactions between task-specific instance queries (ego, agent, and map) and dense Bird’s Eye View (BEV) features while incorporating intent uncertainty in the ego-vehicle’s planning queries. Subsequently, cross-task instance interaction is conducted among task-specific sparse representations with inter-dependencies, where trajectory uncertainty is explicitly modeled in the ego-vehicle’s planning instance. Next, we introduce the hierarchical BEV-instance perception paradigm and the hierarchical intent-trajectory planning mechanism.
@@ 3.1 Hierarchical Perception
$\begin{split}Q_{ego}&=E_{ego}+E_{cmd}+E_{intent}\\ I_{ego}&=Transformer(Q=Q_{ego},\ \ K,V=Adapter_{ego}(F_{bev}))\end{split}$ (1)
@@
Equation (1) models the interaction between planning queries and BEV features $F_{bev}$. CogAD integrates ego embedding $E_{ego}$, high-level driving commands (e.g., lane-keeping/turning) $E_{cmd}$, and intent anchors $E_{intent}$ to initialize the ego-vehicle queries $Q_{ego}$, which are then fed into the EgoTransformer module shown in Figure 2. The intent anchors encode the ego vehicle’s coarse planning space, with implementation details provided in Section 3.2.
@@
CogAD also performs online mapping and dynamic obstacle detection via the interaction between BEV features and map/motion queries, constituting mission-critical modules that are algorithmically coupled with the ego planning task. Online mapping provides topological constraints for planning. CogAD employs multiple learnable spatial interaction queries to decode structured road topology from BEV features through multi-head cross attention, implemented through the MapTransformer module in Figure 2. This interaction process can be formally defined as:
$I_{map}=Transformer(Q=Q_{map},\ \ K,V=Adapter_{map}(F_{bev}))$ (2)
@@
Obstacle detection provides precise spatial localization of surrounding agents, thereby ensuring safety-critical planning. CogAD initiates with interaction queries that extract spatiotemporal information from BEV features via the DetTransformer module in Figure 2, whereby geometric attributes (position, heading angle) and semantic properties (classification confidence) are jointly predicted via a lightweight multi-layer perceptron. The interaction process can be formulated as follows:
$I_{agent}=Transformer(Q=Q_{agent},\ \ K,V=Adapter_{agent}(F_{bev}))$ (3)
Cross-Task Instance Interaction.
Although BEV features facilitate a unified scene understanding, they exhibit limitations in representing the multimodal future trajectories of dynamic agents and ego-vehicle. Additionally, gradient conflicts [6, 21] are likely to arise among the loss functions of multiple tasks, within the BEV space. To address these limitations, CogAD involves inter-instances interactions across different tasks, as shown in the upper right part of Figure 2.
In the planning module, we account for trajectory uncertainty by introducing motion mode embeddings, as elaborated in Section 3.2. Regarding obstacles, we enhance their representations by incorporating dynamic attributes through motion prediction, aiming to avoid potential collisions with the ego-vehicle’s planned trajectory. Similarly, we introduce motion mode embeddings to facilitate multi-modal motion prediction. We align with the philosophy advocated by SparseDrive [47], which emphasizes that motion prediction and planning should account for bidirectional interactions. To this end, the instance queries for ego-vehicle planning and motion prediction are concatenated and then fed into a self-attention (SA) module, thereby facilitating bidirectional information interaction between the two tasks. This design enables the planning task to consider potential future behaviors of surrounding agents, while the motion prediction task also benefits from the planned trajectory of the ego-vehicle. This interaction process can be formulated as follows:
$I_{ego}^{{}^{\prime\prime}},I_{mot}^{{}^{\prime\prime}}=Transformer(Q,K,V=Transformer(Q=[I_{ego}^{{}^{\prime}},I_{mot}^{{}^{\prime}}],K,V=I_{map}))$ (5)
Figure 3: Intent-level and Trajectory-level uncertainty modeling.
3.2 Hierarchical Planning with Uncertainty Modeling
Intent Uncertainty.
Unlike many LLM-based autonomous driving methods [50, 25] that rely on hand-crafted meta-actions to represent intent, CogAD leverages a data-driven approach to obtain intent points. Following practices similar to those in VADv2 [5] and Hydra-MDP [30], we employ K-means clustering to generate trajectory anchors, which serve as intent prototypes. To simplify the implementation, we utilize online K-means for constructing trajectory anchors. Notably, as our focus is on high-level intent discovery rather than fine-grained trajectory clustering, the cluster count is significantly lower than these previous methods [5, 30]. After the hierarchical perception, the ego queries are pooled along the anchor dimension and subsequently processed through the intent head to compute the intent classification loss or derive the selected anchor index, as illustrated in the left part of Figure 3.
Trajectory Uncertainty.
Drawing inspiration from MultiPath++ [52], our hypothesis posits that trajectory uncertainty follows certain intrinsic motion patterns. To capture these motion patterns, we directly employ learnable mode embeddings to represent them. We argue that these motion patterns demonstrate consistency across both planning and motion prediction tasks, given that both tasks fundamentally involve forecasting the future trajectories of agents. Therefore, we utilize shared mode embeddings for both planning and motion prediction tasks. Our experimental results indicate that this strategy not only enhances model performance, but also effectively avoids mode collapse [4, 42, 17, 33]. Finally, based on the anchor selected by the intent head, the trajectory head generates multi-mode trajectories. During training, a winner-takes-all strategy [52, 28] is employed to compute the regression loss and classification loss, as shown on the right part of Figure 3.
3.3 Training and Inference
Training.
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{map}}+\lambda_{2}\mathcal{L}_{\text{det}}+\lambda_{3}\mathcal{L}_{\text{mot}}+\lambda_{4}\mathcal{L}_{\text{plan\_intent}}+\lambda_{5}\mathcal{L}_{\text{plan\_wta}}+\lambda_{6}\mathcal{L}_{\text{plan\_constr}}+\lambda_{7}\mathcal{L}_{\text{k-means}}$ (6)
where $\mathcal{L}_{\text{map}}$ denotes the online mapping loss, $\mathcal{L}_{\text{det}}$ denotes the obstacle detection loss, $\mathcal{L}_{\text{mot}}$ denotes the motion prediction loss, $\mathcal{L}_{\text{plan\_intent}}$ denotes the intent grounding loss, $\mathcal{L}_{\text{plan\_wta}}$ denotes the multi-mode trajectory imitation loss, $\mathcal{L}_{\text{plan\_constr}}$ denotes the overall planning constraints, and the $\mathcal{L}_{\text{k-means}}$ drives the online K-means clustering to generate intent anchors.
Inference.
4 Experiments
Our experiments are conducted on the nuScenes dataset [3] and the Bench2Drive benchmark [22] for open-loop and closed-loop evaluations, respectively. CogAD plans a 3s future ego-trajectory without using any form of ego state or history information as input. More details about dataset and implementation are provided in the supplementary material.
4.1 Main Results
Open-loop Planning Results.
Table 1: Planning performance on nuScenes dataset. We use * to denote the LLM-based method and ‡ to denote taking ego status information as input.
@@ 4.1 Main Results
We compare CogAD against recent state-of-the-art (sota) end-to-end methods, including LLM-based planners. As shown in Table 1, CogAD demonstrates significant advantages in both performance and efficiency. Specifically, CogAD achieves a sota average collision rate of 0.06, substantially outperforming previous methods, including those enhanced with LLM. Meanwhile, CogAD maintains competitively low L2 error and high computational efficiency. To our knowledge, CogAD is the first to achieve 0% collision rate at the 1s planning horizon without relying on ego-state or historical ego information. CogAD demonstrates superior overall performance in terms of L2 error, collision rate, and inference latency. Notably, this is achieved without employing resource-intensive techniques such as the massive parameter counts used by DriveTransformer [23] or the post-selection refinement required by SparseDrive [47]. These results highlight the effectiveness and efficiency of CogAD in real-world autonomous driving scenarios.
Long-tail Scenario Results.
Following TOKEN [49], we construct a long-tail scenario validation set curated from nuScenes for comprehensive evaluation. As shown in Table 2, CogAD achieves significant performance improvements over previous methods. Notably, CogAD establishes new sota results in L2 error for all three long-tail scenarios. In the challenging "3-point turn", a zero-shot case absent from training data, CogAD achieves a 10% reduction in L2 error compared to previous sota result. For the "overtake" scenario, CogAD achieves a competitively low collision rate that nearly matches TOKEN [49]’s performance, while simultaneously outperforming it in average L2 error metric by 0.18m.
Closed-loop Planning Results.
As shown in Table 3, we further evaluate CogAD’s closed-loop performance on Bench2Drive. Compared to methods with comparable model sizes, CogAD achieves sota performance in Driving Score and Success Rate metrics, while maintaining competitive results in Efficiency and Comfortness. Furthermore, although DriveTransformer [23] employs 10$\times$ more parameters, CogAD demonstrates superior Efficiency (42% improvement) and Comfortness (100% better) metrics, highlighting its exceptional balance between model complexity and performance.
@@
Method Closed-loop Metric Latency(ms)
DS $\uparrow$ SR (%) $\uparrow$ Efficiency $\uparrow$ Comfortness $\uparrow$
AD-MLP [59] 18.05 0.00 48.45 22.63 3
UniAD [19] 45.81 16.36 129.21 43.58 663.4
VAD [24] 42.35 15.00 157.94 46.01 278.3
GenAD [61] 44.81 15.91 - - -
MomAD [45] 44.54 16.71 170.21 48.63 -
DriveTransformer [23] 63.46 35.01 100.64 20.78 211.7
CogAD 48.30 24.00 142.00 40.37 121.3
4.2 Ablation Study
Effect of Designs in Hierarchical Perception.
@@ 4.2 Ablation Study
1s 2s 3s Avg. 1s 2s 3s Avg.
- $\surd$ 0.27 0.49 0.81 0.52 0.05 0.11 0.23 0.13
$\surd$ - 0.26 0.45 0.75 0.49 0.03 0.09 0.23 0.12
Hierarchical Perception $\surd$ $\surd$ 0.24 0.45 0.74 0.48 0.00 0.02 0.15 0.06
@@
1s 2s 3s Avg. 1s 2s 3s Avg.
- $\surd$ 0.26 0.49 0.82 0.52 0.05 0.11 0.32 0.16
$\surd$ - 0.22 0.41 0.69 0.44 0.09 0.13 0.27 0.16
Hierarchical Planning $\surd$ $\surd$ 0.24 0.45 0.74 0.48 0.00 0.02 0.15 0.06
To validate the effectiveness of the hierarchical perception paradigm, we conduct two ablation studies: one retaining only the ego queries’ interaction with BEV features, and the other preserving solely the ego-to-instance queries interaction. As shown in Table 4, interaction with BEV features benefits both the L2 error and collision rate, while instance-level interaction predominantly reduces the collision rate. These results collectively validate the effectiveness of our hierarchical perception paradigm.
Effect of Designs in Hierarchical Planning.
@@
BEV Adapter Mode Sharing Ego-Motion Interaction 1s 2s 3s Avg. 1s 2s 3s Avg.
- $\surd$ $\surd$ 0.25 0.47 0.80 0.51 0.47 0.55 0.78 0.60
$\surd$ - $\surd$ 0.22 0.41 0.68 0.44 0.05 0.15 0.18 0.13
$\surd$ $\surd$ - 0.25 0.46 0.77 0.49 0.05 0.10 0.25 0.13
$\surd$ $\surd$ $\surd$ 0.24 0.45 0.74 0.48 0.00 0.02 0.15 0.06
Table 4 demonstrates the effectiveness of our hierarchical planning design. The absence of intent uncertainty degrades CogAD’s planning module to a conventional multi-mode ego-trajectory planning method, resulting in a significantly higher collision rate. Besides, preserving only intent uncertainty while omitting trajectory uncertainty also yields suboptimal results, as this configuration eliminates the model’s capacity to generate refined planning trajectories.

### v3

# Cognitive-Hierarchy Guided End-to-End Planning for Autonomous Driving

@@ 1 Introduction
In the perception pattern, we argue that the perception of human drivers is hierarchical. Cognitive psychology supports the pre-attentive two-stage theory of human visual perception [27], which posits that individuals first perform global scanning of the scene to identify contextual or spatial cues, followed by selective attention allocation to critical local objects for detailed analysis. This mechanism aligns with neurobiological evidence [13] demonstrating that early visual areas prioritize scene-wide features, while higher-order regions refine task-specific processing through context-dependent modulation. In terms of driving environmental perception, human drivers typically initiate their environmental perception by establishing holistic scene comprehension of the traffic scenario, encompassing all contextual elements as illustrated in Fig.˜1 (c), followed by selective attention prioritization towards critical dynamic objects such as proximal vehicles, pedestrians, and lane markings, as illustrated in Fig.˜1 (d).
In the planning pattern, we argue that the planning of human drivers is hierarchical. The BDI (Belief-Desire-Intention) cognitive model [2, 41] decomposes human action planning into hierarchically organized processes, distinguishing between high-level intentional states (e.g., goal setting) and low-level behavioral executions (e.g., motor sequence implementation). In terms of driving planning, human drivers typically follow a hierarchical planning process: intentional planning at the higher level (e.g., lane change) followed by trajectory planning at the lower level (e.g., generating a specific trajectory). For instance, drivers first establish a global intent, which then constrains the generation of local trajectories, as illustrated in Fig.˜1 (b).
Figure 1: Human drivers scan surrounds (c) before focusing on key objects (d), and plan hierarchically from intent to trajectory (b).
From the perspective of hierarchical cognitive psychology, recent methods like [56, 25, 19] fail to incorporate hierarchical perception, while others like [47, 23, 33] lack hierarchical planning, as detailed in Sec.˜2. These analyses highlight a critical gap in aligning E2E-AD models with human drivers’ cognitive principles. To bridge this gap, we propose CogAD, an end-to-end autonomous driving model that emulates human hierarchical cognition. For hierarchical perception, we design a sequential interaction mechanism where the ego-vehicle first processes global BEV features to capture environmental context, then focuses on critical instance-level elements. For hierarchical planning, CogAD first plans high-level driving intents, then generates corresponding low-level trajectories, mirroring human multi-level planning.
@@ 2 Related Work
Significant progress has been made in end-to-end autonomous driving (E2E-AD) approaches, pioneered by prior works such as [1, 10, 38].
@@ 3 Method
3.1 Overview
As depicted in Fig.˜2, the proposed CogAD implements a unified multi-task architecture that integrates perception, prediction, and planning in a coordinated manner. To comprehensively capture environmental information, we design a hierarchical feature interaction framework. First, we establish interactions between task-specific instance queries (ego, agent, and map) and dense Bird’s Eye View (BEV) features while incorporating intent uncertainty in the ego-vehicle’s planning queries. Subsequently, cross-task instance interaction is conducted among task-specific sparse representations with inter-dependencies, where trajectory uncertainty is explicitly modeled in the ego-vehicle’s planning instance. Next, we introduce the hierarchical BEV-instance perception paradigm and the hierarchical intent-trajectory planning mechanism.
@@ 3.2 Hierarchical Perception
$\begin{split}Q_{ego}&=E_{ego}+E_{cmd}+E_{intent}\\ I_{ego}&=Transformer(Q=Q_{ego},\\ &\hskip 65.44142ptK,V=Adapter_{ego}(F_{bev}))\end{split}$ (1)
@@
Eq.˜1 models the interaction between planning queries and BEV features $F_{bev}$. $Transformer$ represents the standard transformer module [53]. CogAD integrates ego embedding $E_{ego}$, high-level driving commands (e.g., lane-keeping/turning) $E_{cmd}$, and intent anchors $E_{intent}$ to initialize the ego-vehicle queries $Q_{ego}$, which are then fed into the Ego Transformer module shown in Fig.˜2. The intent anchors encode the ego vehicle’s coarse planning space, with implementation details provided in Sec.˜3.3.
@@
CogAD also performs online mapping and dynamic obstacle detection via the interaction between BEV features and map/motion queries, constituting mission-critical modules that are algorithmically coupled with the ego planning task. Online mapping provides topological constraints for planning. CogAD employs multiple learnable spatial interaction queries to decode structured road topology from BEV features through multi-head cross attention, implemented through the MapTransformer module in Fig.˜2. This interaction process can be formally defined as:
$\begin{split}I_{map}&=Transformer(Q=Q_{map},\\ &\hskip 65.44142ptK,V=Adapter_{map}(F_{bev}))\end{split}$ (2)
@@
Obstacle detection provides precise spatial localization of surrounding agents, thereby ensuring safety-critical planning. CogAD initiates with interaction queries that extract spatiotemporal information from BEV features via the DetTransformer module in Fig.˜2, whereby geometric attributes (position, heading angle) and semantic properties (classification confidence) are jointly predicted via a lightweight multi-layer perceptron. The interaction process can be formulated as follows:
$\begin{split}I_{agent}&=Transformer(Q=Q_{agent},\\ &\hskip 65.44142ptK,V=Adapter_{agent}(F_{bev}))\end{split}$ (3)
Table 1: Planning performance on nuScenes dataset. We use * for LLM-based method.
@@
Cross-Task Instance Interaction.
Although BEV features facilitate a unified scene understanding, they exhibit limitations in representing the multimodal future trajectories of dynamic agents and ego-vehicle. Additionally, gradient conflicts [6, 21] are likely to arise among the loss functions of multiple tasks, within the BEV space. To address these limitations, CogAD involves inter-instances interactions across different tasks, as shown in the upper right part of Fig.˜2.
In the planning module, we account for trajectory uncertainty by introducing motion mode embeddings, as elaborated in Sec.˜3.3. Regarding obstacles, we enhance their representations by incorporating dynamic attributes through motion prediction, aiming to avoid potential collisions with the ego-vehicle’s planned trajectory. Similarly, we introduce motion mode embeddings to facilitate multi-modal motion prediction. We align with the philosophy advocated by SparseDrive [47], which emphasizes that motion prediction and planning should account for bidirectional interactions. To this end, the instance queries for ego-vehicle planning and motion prediction are concatenated and then fed into a self-attention (SA) module, thereby facilitating bidirectional information interaction between the two tasks. This design enables the planning task to consider potential future behaviors of surrounding agents, while the motion prediction task also benefits from the planned trajectory of the ego-vehicle. This interaction process can be formulated as follows:
$\begin{split}I_{ego}^{{}^{\prime\prime}},I_{mot}^{{}^{\prime\prime}}=Transformer(Q,K,V=\hskip 51.21504pt\\ Transformer(Q=[I_{ego}^{{}^{\prime}},I_{mot}^{{}^{\prime}}],K,V=I_{map}))\end{split}$ (5)
Figure 3: Intent-level and Trajectory-level uncertainty modeling.
3.3 Hierarchical Planning
Intent Uncertainty.
Unlike many LLM-based autonomous driving methods [50, 24] that rely on hand-crafted meta-actions to represent intent, CogAD leverages a data-driven approach to obtain intent points. Following practices similar to those in VADv2 [5] and Hydra-MDP [30], we employ K-means clustering to generate trajectory anchors, which serve as intent prototypes. To simplify the implementation, we utilize online K-means for constructing trajectory anchors. Notably, as our focus is on high-level intent discovery rather than fine-grained trajectory clustering, the cluster count is significantly lower than these previous methods [5, 30]. After the hierarchical perception, the ego queries are pooled along the anchor dimension and subsequently processed through the intent head to compute the intent classification loss or derive the selected anchor index, as illustrated in the left part of Fig.˜3.
Trajectory Uncertainty.
Drawing inspiration from MultiPath++ [52], our hypothesis posits that trajectory uncertainty follows some intrinsic motion patterns. To capture these motion patterns, we directly employ learnable mode embeddings to represent them. We argue that these motion patterns demonstrate consistency across both planning and motion prediction tasks, given that both tasks fundamentally involve forecasting the future trajectories of agents. Therefore, we utilize shared mode embeddings for both planning and motion prediction tasks. Our experimental results indicate that this strategy not only enhances model performance, but also effectively avoids mode collapse [4, 42, 17, 33]. Finally, based on the anchor selected by the intent head, the trajectory head generates multi-mode trajectories. During training, a winner-takes-all strategy [52, 28] is employed to compute the regression loss and classification loss, as shown on the right part of Fig.˜3.
3.4 Training and Inference
Training.
$\begin{split}\mathcal{L}=&\lambda_{1}\mathcal{L}_{\text{map}}+\lambda_{2}\mathcal{L}_{\text{det}}+\lambda_{3}\mathcal{L}_{\text{mot}}+\lambda_{4}\mathcal{L}_{\text{plan\_intent}}+\\ &\lambda_{5}\mathcal{L}_{\text{plan\_wta}}+\lambda_{6}\mathcal{L}_{\text{plan\_constr}}+\lambda_{7}\mathcal{L}_{\text{k-means}}\end{split}$ (6)
where $\mathcal{L}_{\text{map}}$ denotes the online mapping loss, $\mathcal{L}_{\text{det}}$ denotes the obstacle detection loss, $\mathcal{L}_{\text{mot}}$ denotes the motion prediction loss, $\mathcal{L}_{\text{plan\_intent}}$ denotes the intent grounding loss, $\mathcal{L}_{\text{plan\_wta}}$ denotes the multi-mode trajectory imitation loss, $\mathcal{L}_{\text{plan\_constr}}$ denotes the overall planning constraints, and the $\mathcal{L}_{\text{k-means}}$ drives the online K-means clustering to generate intent anchors. Eq.˜6 is optimized by joint training all CogAD components, with the loss weighting parameters set to $\lambda_{1}=\lambda_{2}=2.0$, $\lambda_{3}=0.2$, and $\lambda_{4}=\lambda_{5}=\lambda_{6}=\lambda_{7}=1.0$.
@@ 3.4 Training and Inference
Inference.
4 Experiments
Our experiments are conducted on the nuScenes dataset [3] and the Bench2Drive benchmark [22] for open-loop and closed-loop evaluations, respectively. CogAD intentionally excludes any form of ego state or historical trajectory as input to avoid overfitting [32, 60]. For fairness, we compare only against methods that do not utilize this information and adhere to the same training protocol of 60 epochs on nuScenes and 6 epochs on Bench2Drive. More details about dataset and implementation details are provided in the supplementary material.
@@ 4 Experiments
Method DS $\uparrow$ SR (%) $\uparrow$ Latency(ms)
UniAD [19] 45.81 16.36 663.4
VAD [25] 42.35 15.00 278.3
GenAD [62] 44.81 15.91 -
MomAD [45] 44.54 16.71 -
CogAD 48.30 24.00 121.3
4.1 Main Results
@@
[…]
@@
[…]
@@
[…]
Open-loop Planning Results.
We compare CogAD against recent state-of-the-art (sota) end-to-end methods, including LLM-based planners. As shown in Tab.˜1, CogAD demonstrates significant advantages in both performance and efficiency. Specifically, CogAD achieves a sota average collision rate of 0.06, substantially outperforming previous methods, including those enhanced with LLM. Meanwhile, CogAD maintains competitively low L2 error and high computational efficiency. To our knowledge, CogAD is the first to achieve 0% collision rate at the 1s planning horizon without relying on ego-state or historical ego information. CogAD demonstrates superior overall performance in terms of L2 error, collision rate, and inference latency. Notably, this is achieved without employing resource-intensive techniques, such as the massive parameter counts or the post-selection refinement. These results highlight the effectiveness and efficiency of CogAD in real-world autonomous driving scenarios.
Long-tail Scenario Results.
Following TOKEN [49], we construct a long-tail scenario validation set curated from nuScenes for comprehensive evaluation. As shown in Tab.˜2, CogAD achieves significant performance improvements over previous methods. Notably, CogAD establishes new sota results in L2 error for all three long-tail scenarios. In the challenging "3-point turn", a zero-shot case absent from training data, CogAD achieves a 10% reduction in L2 error compared to previous sota result. For the "overtake" scenario, CogAD achieves a competitively low collision rate that nearly matches TOKEN [49]’s performance, while simultaneously outperforming it in average L2 error metric by 0.18m.
Closed-loop Planning Results.
As shown in Tab.˜3, the closed-loop evaluation of CogAD on the challenging Bench2Drive benchmark demonstrates its superior performance. Compared to other methods, CogAD achieves state-of-the-art performance in Driving Score and Success Rate, alongside competitive latency.
4.2 Ablation Study
Effect of Designs in Hierarchical Perception.
To validate the effectiveness of the hierarchical perception paradigm, we conduct two ablation studies: one retaining only the ego queries’ interaction with BEV features, and the other preserving solely the ego-to-instance queries interaction. As shown in Tab.˜4, interaction with BEV features benefits both the L2 error and collision rate, while instance-level interaction predominantly reduces the collision rate. These results collectively validate the effectiveness of our hierarchical perception paradigm.
Effect of Designs in Hierarchical Planning.
Tab.˜4 demonstrates the effectiveness of our hierarchical planning design. The absence of intent uncertainty degrades CogAD’s planning module to a conventional multi-mode ego-trajectory planning method, resulting in a significantly higher collision rate. Besides, preserving only intent uncertainty while omitting trajectory uncertainty also yields suboptimal results, as this configuration eliminates the model’s capacity to generate refined planning trajectories. Our visualization analysis (some videos in the Supplementary Material) confirms that hierarchical planning delivers critical advantages in collision avoidance, enabling maneuvers like deceleration for yielding and lane-changing to bypass obstacles. This explains the results in Tab.˜4: Collision Rate decreases by 62.5% versus non-hierarchical planning counterparts, while L2 Error remains statistically comparable (showing minimal variation).
Figure 4: Qualitative results of CogAD on nuScenes. The 2nd column shows the Top-3 multi-mode trajectories of the highest-probability intent, with colors of red, orange, and yellow, respectively. The 3rd column displays the Top-10 multi-intent trajectories with the highest-probability mode. GT trajectory is drawn in green.